OpenStudy (anonymous):

Graph the inequality three times n minus four is less than or equal to twenty-three.

OpenStudy (zpupster):

hello lets set it up: n=x ( the limit of the graphing software so 3x-4<= to 23 just like the last problem lets isolate x do you know how??

OpenStudy (anonymous):

umm you get it by itself right so to do that we have to add by 4 right?

OpenStudy (zpupster):

yes that is right lets add 4 to both sides

OpenStudy (anonymous):

ok

OpenStudy (zpupster):

3x-4+4<=to 23+4

OpenStudy (anonymous):

23+4=27

OpenStudy (zpupster):

3x<=to 27 now what??

OpenStudy (anonymous):

27/3 right which equals 9

OpenStudy (zpupster):

yes and the 3 cancel on the left so x<= to 9

OpenStudy (anonymous):

thanks

OpenStudy (zpupster):

here the solution in graph

OpenStudy (anonymous):

ok so like this